OpenStudy (anonymous):

What is the simplified mixed number form of 22/10? A 1 12/10 B 1 1/5 C 2 1/5

OpenStudy (anonymous):

\[1\frac{ 12 }{ 10 } ,,,,,,,,,,, 1\frac{ 1 }{ 5 } ,,,,,,OR,,,,, 2\frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

Ok thank you :)

OpenStudy (anonymous):

sorry just realised your question wants the most simplified and thus C is the most simplified

OpenStudy (anonymous):

Oh ok lol thanks^^